OpenStudy (anonymous):

A equilateral triangle has a height of 30 yards. What is the area and perimeter?

OpenStudy (anonymous):

the perimeter is the easy part, you just add the three sides that are 30 yards each for a total of 90 yards. area of a triangle is equal to (1/2)*base*height. your base is 30, so 1/2 that would be 15 and your height can be found using the Pythagorean theorem. 15^2+x^=30^2 30^2-15^2=x^2 x=sqrt(675) ~ 25.98 then you have your base: 15, your height ~25.98, so you just need to multiply those together for a product of approximately 389.711